I will rate. thank you

N2(g)+ 3 H2(g)= 2 NH3

Find delta G of reaction when mixing 4 mol of nitrogen, 12 moles of of h2 and 4 mol nh3 (pressure = 1 bar) temp = 50c
Answer #1

Nz1g) + 3H₂(g) + 2NH319) temp=T=500 = 50+273=323k For this reaction equilibrium expression (constant) C K. - [NH² [N2] [11] [Here, volume is not given means volume I is constant thats why concentration is in mole. From ① & Ke = (4) ² - 16 (4) (123 -for equilibrium reaction, obso - . o=0G + RT enke 7. AG =-RT anke T= 323k, R= 8.3145 J/kmol AG = – 8.3145 X 323KX 2.303 xlookG=-6184.8988 logke = - 6/84-8988 10g (2.3148X103) = - 6184.8988X(-2.6354) = 16299.6823 Ilmol 8 = 16.2996 KJ/mol 3